OpenStudy (anonymous):

Use the binomial expression (p + q)n to calculate a binomial distribution with n = 5 and p = 0.3.

OpenStudy (anonymous):

(p+q)n =(0.3+q)5 =1.5+5q distribute the 5 into what is in the ( ) 5*0.3=1.5 5*q=5q =1.5+5q

OpenStudy (anonymous):

This was for a test that I had to retake. I put that exact answer and she told me it was wrong

OpenStudy (anonymous):

(p + q)^5 = p^5 + (5p^4)(q) + (10p^3)(q^2) + (10 p^2)(q^3) + 5p(q^4) + q^5 If p = 0.3, then q = 1 - p = 1 - 0.3 = 0.7.

OpenStudy (anonymous):

Ask her how it is wrong

OpenStudy (anonymous):

Because the 5 she told me what useless. Like they throw it in to confuse you. But I figured it out. Thanks guys(:

OpenStudy (perl):

Let X = number of successes, you have a binomial distribution. and p is the probability of success, q is the probability of failure. X can equal 0,1,2,3,4,5 P(X = x) = nCx p^x q^(n-x)